OpenStudy (sierra!):

WILL MEDAL AND FAN!!!!!!!!!!!! Which mixed number is equal to 7.6?

OpenStudy (sierra!):

A. 1/4 = 0.25 _ B. 1/3 =0.3 C. 1/2 = 0.5 _ D. 2/3 = 0.6

OpenStudy (anonymous):

7.6 The 7 stays as a 7, as it is a mixed number. 0.6 is simply 6/10, which simplifies to 3/5. So it ends up as 7 and 3/5

OpenStudy (sierra!):

huh @Balvasor

OpenStudy (sierra!):

@nincompoop

OpenStudy (sierra!):

@iambatman @Compassionate

OpenStudy (sierra!):

@mathstudent55 @Luigi0210 @pooja195

OpenStudy (sierra!):

@sleepyjess

OpenStudy (sierra!):

@zepdrix

OpenStudy (sierra!):

@bibby

OpenStudy (sierra!):

@ganeshie8 @Nnesha